OpenStudy (anonymous):

Explain how you would find the product of 20^12 and 20^8. Include either the exponent rule. Give your answer in exponential form

OpenStudy (the_fizicx99):

Well the bases are the same so just add 12 + 8 20^20

OpenStudy (wolf1728):

20 ^ 12 • 20 ^ 8 Basically, when multiplying numbers whose bases are the same number. raised to different powers, the result is 20^x • 20^y = 20^(x + y) So 20 ^12 • 20 ^ 8 = 20 ^ 20